Add the following. 
a – b + ab, b – c + bc, c – a + ac

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

(a - b + ab) + (b - c + bc) + (c - a + ac)
= a - b + ab + b - c + bc + c - a + ac
= a - a +b–b +c–c + ab + bc + ca
=0 + 0 + 0 + ab + bc + ca
= ab + bc + ca

Thanks